OpenStudy (anonymous):

can som one hlp me plzz Solve using substitution: y − 2x − 1 = 0 -4y = 6 − 8x Answer (1/2, 2) (3, 1) No solution Infinitely many solutions

OpenStudy (anonymous):

cn sme one hlp

OpenStudy (anonymous):

can yu hlp Kaylei104

OpenStudy (anonymous):

Im going to start with the second problem. -4y=6-8x divide both sides by -4 to get y=2x-3/2 then take the first problem and plug 2x-3/2 in for y so our first equation should be (2x-3/2)-2x-1=o

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

once we have 2x-3/2-2x-1=0 we are going to solve for x. 2x-2x is 0 and -3/2-1 is -5/2 we end up with -5/2=0 which is impossible so there is no solution

OpenStudy (anonymous):

do you understand how I found the answer?

OpenStudy (anonymous):

no